iuytyuioiuytyuiop:

Find the equation of the line with slope m=−13 and passing through the point (1,13).

iuytyuioiuytyuiop:

@Angle

Angle:

We have something called ` point-slope form ` ` y = m(x - a) + b ` when m is the slope and the point (a, b) is a point on the line

iuytyuioiuytyuiop:

y=-1/3(x-1) + 1/3

Angle:

you got it :)

iuytyuioiuytyuiop:

do I multiply by 5?

Angle:

nope, we only did that last time because the standard form doesn't like fractions

iuytyuioiuytyuiop:

but 1/3 and 1/3 are fractions

Angle:

yes, but what we have is point-slope form only the standard form doesn't like fractions does your question want your equation in standard form?

iuytyuioiuytyuiop:

Yep it has to be in standard form

Angle:

ohhh then yes, we want to multiply by the common denominator so multiply by 3 this time

iuytyuioiuytyuiop:

y=-1(x-3)+1

Angle:

awesome! then if we want standard form, we first want to distribute that -1 into the parentheses

iuytyuioiuytyuiop:

y=(-1-3)+1

Angle:

close -1(x - 3) = -x + 3

iuytyuioiuytyuiop:

y=(-x-3)+1

Angle:

negative times negative becomes a positive

iuytyuioiuytyuiop:

y=(x-3)+1

Angle:

y = -x + 3 + 1

Angle:

which simplifies to y = -x + 4 then move the x to the other side

iuytyuioiuytyuiop:

y=-4x

iuytyuioiuytyuiop:

y=4x

Angle:

x + y = 4

iuytyuioiuytyuiop:

it's x + y = 4?!!???????????????!!!!!!

Angle:

mhm

iuytyuioiuytyuiop:

thanks for helping me I'm grateful for your help

iuytyuioiuytyuiop:

@Angle its incorrect

Angle:

ah I forgot to multiply the y by 3 as well

iuytyuioiuytyuiop:

LOl

Angle:

x + 3y = 4

iuytyuioiuytyuiop:

thanks

iuytyuioiuytyuiop:

it's x+3y=2

Angle:

weird

iuytyuioiuytyuiop:

lol